A treasury inflation protected security has a starting par value of 1000, and a coupon rate of 5% paid annually. At the end of two years, inflation was reported at 2% and 4% respectively. What will be the coupon paid on the bond after these two years?
Answer #1

TIPS Treasury inflation-protected security adjusts themselves to the inflation by adjusting its par value.

The adjusted par value of the Bond = par value at starting x(1+ inflation rate).

At the end of year 1 = 1000(1+0.02)
=1020

At the end of year 2 = 1020(1+0.04)
=1060.8

the coupon paid on the bond after these two years = latest adjusted par value of the bond x coupon rate.

=1060.8×5%
= $53.04